OpenStudy (anonymous):

how do you find the y intercept and slope when y=8.5x-8?

OpenStudy (anonymous):

This is in the form y=mx+c. Remember that m is the gradient or slope, and c is the y intercept. Can you work it out now? :)

OpenStudy (anonymous):

so is this a LinReg ax+b test on my calculator?

OpenStudy (anonymous):

I'm sorry, I have no idea what that is :S But if you just need the slope and y intercept, you don't need a calculator!

OpenStudy (anonymous):

ohh ok, so I still can't figure it out

OpenStudy (anonymous):

Hm, what do you need to do?

OpenStudy (anonymous):

can you step me through it?

OpenStudy (anonymous):

What do you need?

OpenStudy (anonymous):

ok so if y=Mx+C ...does that mean I do 8.5x+8?

OpenStudy (anonymous):

You do that when you want to find out the value of y or x. If you want the slope and y intercept you don't need to do this. y=mx+c m=slope c=y intercept y=8.5x-8 m=8.5 c=-8 So the slope is 8.5, and the y intercept is -8 :)

OpenStudy (anonymous):

ohh ok got it! THX!!!

OpenStudy (anonymous):

Awesome! :) No problem!